# Approximately Dominating Sets in Elections

### v2

@@ 1 Introduction
However, as [ELS11, ELS15] note, one strange feature of Condorcet winning sets is that they can be quite far from dominating sets. For example, a 1-winning set may exclude a candidate that would beat each member in the committee by a wide margin. To see this, consider an election where voters’ first choices are distinct but their second choices are unanimous. The first choices would be a 1-winning set, but for any of these candidates, nearly all voters would prefer the second choice (which is in fact a Condorcet winner in the strongest sense). [ELS11, ELS15] also show empirically that the Condorcet winning sets and dominating sets in random elections are usually very different.
@@
On the other hand, the recent work on $\theta$-winning sets can be used to obtain positive results for small $\alpha$. It is not hard to see that while any $\alpha$-dominating set is always a $\theta$-winning set, a $\theta$-winning set of size $k$ is at the very least a $\frac{\theta}{k}$-dominating set. With this, [CLR+25] imply that for for $\alpha\leq 0.101$, $k_{\alpha}=2$. Though they give stronger guarantees on $\theta$-winning sets for larger committee sizes, they do not yield stronger guarantees for $\alpha$-dominating sets (due to the factor $k$ dilution).
@@ 1.1 Approximating Maximal Lotteries
With this game-theoretic correspondence in mind, what Theorem 1 shows is that we can approximate maximal lotteries with constant support. More precisely, we can construct a mixed strategy with support size $O(\frac{1}{\varepsilon^{2}})$ that guarantees expected payoff at least $-\varepsilon$ in the aforementioned game.
@@
All together, these connections make it somewhat surprising that maximal lotteries can be approximated with constant support size, without the $\log m$ dependence one might expect. It is also fortuitous that the construction and proof are simple, and it would be interesting to study whether similar techniques can extend to other specific approximate Nash equilibria of interest. Moreover, given the vast variety of research developing and using maximal lotteries, approximate maximal lotteries may be of independent interest.
@@ 1.2 Further Related work
Nguyen, Song, and Lin [NSL25] very recently posted a new preprint, improving [CLR+25]’s bounds on $\theta$-winning sets for all committee sizes $k\geq 2$. Though the work is not yet on a public repository, it is currently available through the first author’s webpage.
@@ 2 Preliminaries and Notation
We use $D$ (and variants) to denote distributions over candidates, and $x\sim D$ to say that $x$ is sampled from $D$. With a set $S$ we also write $x\sim S$ to say that $x$ is sampled from the uniform distribution over $S$. (Specifically, $x\sim[0,1]$ denotes sampling a uniform real number between 00 and $1$, and $v\sim V$ denotes sampling a uniformly random voter.)
@@ 3 Proof of Theorem 1
Before diving into the proof, we first provide some high-level intuition for the analysis, particularly for why it is possible to avoid a log factor blow up. Our goal is to construct a distribution with small support that approximates a maximal lottery. This distribution will simply be the empirical distribution of a number of samples from the maximal lottery. To show that the empirical distribution is not too different from a maximal lottery, we want to argue that for each candidate $a$, the average voter’s rank of $a$ with respect to the maximal lottery and with respect to the empirical distribution are not too different. We can make such an argument using concentration inequalities, but if we need to union bound over all $m$ candidates, we will incur a $\log m$ blow up in the number of samples. What saves us is that each voter’s ranks of the candidates always follow the linear order of that voter’s preferences, so if one candidate’s rank changes by only slightly, it limits other candidates with close ranks from changing significantly. Heuristically, if we want to show that all ranks change by at most $O(\varepsilon)$, it is really sufficient to show this for $O(1/\varepsilon)$ evenly distributed ranks, since this will extend to all ranks that are sandwiched in between. This argument would replace the $\log m$ with a $\log\frac{1}{\varepsilon}$.
@@
We note that, along the lines of the earlier heuristic argument, one can also prove a weaker bound of $\delta(k)=O\Big{(}\sqrt{\tfrac{\log k}{k}}\Big{)}$ with less heavy machinery. We can discretize the interval $[0,1]$ into grid points $r=\frac{1}{k},\frac{2}{k},\dots,\frac{k}{k}$, and argue via Hoeffding’s inequality that with high probability, $\widehat{F}(r)$ is within $O\Big{(}\sqrt{\tfrac{\log k}{k}}\Big{)}$ of $r$ for all of these points. Due to the fine discretization and monotonicity of $\widehat{F}(\cdot)$, this extends to all $r\in[0,1]$ with at most an additional $\frac{1}{k}$ error.
@@
$\max_{a\in C}\mathop{\mathbb{E}}_{b\sim D}\big{[}\tfrac{1}{n}|a\succ b]\big{]}\leq\frac{1}{2}+\varepsilon.$
@@
Now suppose we sample $b_{1},\dots,b_{k}\sim D_{\textnormal{ML}}$, where $k$ will eventually be $(1+o(1))\frac{\pi}{8\varepsilon^{2}}$. We claim that with positive probability, the empirical distribution $\widehat{D}$ which samples uniformly from $b_{1},\dots,b_{k}$ satisfies the property required by the theorem. The intuition is to argue that in expectation, for each candidate $a$ and the average voter $v$, $\operatorname{rank}_{v}(a;\widehat{D})$ is not too much larger than $\operatorname{rank}_{v}(a;D_{\textnormal{ML}})$, and so $\widehat{D}$ will satisfy an approximate version of Eq. 2.
From the perspective of each voter $v$, we can view samples from any distribution $D$ over candidates in the following way. We split the interval $[0,1]$ into disjoint segments corresponding to candidates, and the length of the segment for each candidate $a$ is the probability mass of $a$ in $D$. The segments are ordered by $v$’s preference, with more preferred candidates closer to $1$ and less preferred closer to 00. Note that $\operatorname{rank}_{v}(a;D)$ is precisely the minimum of the segment corresponding to $a$.
@@
$\mathop{\mathbb{E}}_{b\sim\widehat{D}}[\tfrac{1}{n}|a\succ b]]$ $=\frac{1}{n}\sum_{v\in V}\operatorname{rank}_{v}(a;\widehat{D})$
@@
Evaluating these integrals for $k=2,3,4$, we can get that $\delta(2)=\frac{3}{8},\delta(3)=\frac{17}{54},\delta(4)=\frac{71}{256}$, which shows that $k_{\alpha}=2$ for $\alpha\leq\frac{1}{8}=0.125$, $k_{\alpha}\leq 3$ for $\alpha\leq\frac{5}{27}\approx 0.185$, and $k_{\alpha}\leq 4$ for $\alpha\leq\frac{57}{256}\approx 0.222$.
@@ Appendix A A note on lower bounds
We will argue that a random subset $B$ of $[t]\setminus A$ of size $b$ satisfies this property with positive probability. For each $i$, the probability that $|A_{i}\cap B|>\frac{b}{2}$ is at most (very loosely)

### v3

@@ 1 Introduction
However, as [ELS11, ELS15] note, one strange feature of Condorcet winning sets is that they can be quite far from dominating sets. For example, a 1-winning set may exclude a candidate that would beat each member of the committee by a wide margin. To see this, consider an election where voters’ first choices are distinct but their second choices are unanimous. The first choices would form a 1-winning set, but for any of these candidates, nearly all voters would prefer the second choice (which is in fact a Condorcet winner in the strongest sense). [ELS11, ELS15] also show empirically that the Condorcet winning sets and dominating sets in random elections are usually very different.
@@
On the other hand, the recent work on $\theta$-winning sets can be used to obtain positive results for small $\alpha$. It is not hard to see that while any $\alpha$-dominating set is always an $\alpha$-winning set, a $\theta$-winning set of size $k$ is at the very least a $\frac{\theta}{k}$-dominating set. With this, [CLR+25] imply that for for $\alpha\leq 0.101$, $k_{\alpha}=2$. Though they give stronger guarantees on $\theta$-winning sets for larger committee sizes, they do not yield stronger guarantees for $\alpha$-dominating sets (due to the factor $k$ dilution).
@@ 1.1 Approximating Maximal Lotteries
With this game-theoretic correspondence in mind, what Theorem 1 shows is that we can approximate maximal lotteries with constant support. More precisely, we can construct a mixed strategy with support size $O(\frac{1}{\varepsilon^{2}})$ that guarantees expected payoff at least $\varepsilon$ in the aforementioned game.
@@
All together, these connections make it somewhat surprising that maximal lotteries can be approximated with constant support size, without the $\log m$ dependence one might expect (where $m$ is the number of candidates). It is also fortuitous that the construction and proof are simple, and it would be interesting to study whether similar techniques can extend to other specific approximate Nash equilibria of interest. Moreover, given the vast variety of research developing and using maximal lotteries, approximate maximal lotteries may be of independent interest.
@@ 1.2 Further Related work
Nguyen, Song, and Lin [NSL25] very recently posted a new preprint, improving [CLR+25]’s bounds on $\theta$-winning sets for all committee sizes $k\geq 2$.
@@
Processes for approximating maximal lotteries.
Brandl and Brandt [BB24] propose a simple and elegant scheme that approximates maximal lotteries with only lightweight preference elicitation. They imagine an urn filled with many balls, each arbitrarily labeled with a candidate. In every round two balls are drawn uniformly at random, and a randomly selected voter is compares the corresponding candidates. The ball representing the losing candidate is replaced with a copy of the winning one. To maintain ergodicity, there is also a small probability that a random ball is randomly relabeled. [BB24] show that with sufficiently many balls, the distribution of labels converges exponentially fast to a maximal lottery.
@@ 2 Preliminaries and Notation
We use $D$ (and variants) to denote distributions over candidates, and $x\sim D$ to say that $x$ is sampled from $D$. With a set $S$ we also write $x\sim S$ to say that $x$ is sampled from the uniform distribution over $S$. (Specifically, $x\sim[0,1]$ denotes sampling a uniform real number between 0 and $1$, and $v\sim V$ denotes sampling a uniformly random voter.)
@@ 3 Proof of Theorem 1
Before diving into the proof, we first provide some high-level intuition for the analysis, particularly for why it is possible to avoid a log factor blow up. Our goal is to construct a distribution with small support that approximates a maximal lottery. This distribution will simply be the empirical distribution of a number of samples from the maximal lottery. To show that the empirical distribution is not too different from a maximal lottery, we want to argue that for each candidate $a$, the average voter’s rank of $a$ with respect to the maximal lottery and with respect to the empirical distribution are not too different. We can make such an argument using concentration inequalities, but if we need to union bound over all $m$ candidates, we will incur a $\log m$ blow up in the number of samples. What saves us is that each voter’s ranks of the candidates always follow the linear order of that voter’s preferences, so if one candidate’s rank changes only slightly, it limits other candidates with close ranks from changing significantly. Heuristically, if we want to show that all ranks change by at most $O(\varepsilon)$, it is really sufficient to show this for $O(1/\varepsilon)$ evenly distributed ranks, since this will extend to all ranks that are sandwiched in between. This argument would replace the $\log m$ with a $\log\frac{1}{\varepsilon}$.
@@
[…]
@@
$\max_{a\in C}\mathop{\mathbb{E}}_{b\sim D}\big[\tfrac{1}{n}|a\succ b|\big]\leq\frac{1}{2}+\varepsilon.$
@@
[…]
From the perspective of each voter $v$, we can view samples from any distribution $D$ over candidates in the following way. We split the interval $[0,1]$ into disjoint segments corresponding to candidates, and the length of the segment for each candidate $a$ is the probability mass of $a$ in $D$. The segments are ordered by $v$’s preference, with more preferred candidates closer to $1$ and less preferred closer to 0. Note that $\operatorname{rank}_{v}(a;D)$ is precisely the minimum of the segment corresponding to $a$.
@@
[…]
@@
Evaluating these integrals for $k=2,3,4$, we can get that $\delta(2)=\frac{3}{8},\delta(3)=\frac{17}{54},\delta(4)=\frac{71}{256}$, which shows that $k_{\alpha}=2$ for $\alpha\leq\frac{1}{8}=0.125$, $k_{\alpha}\leq 3$ for $\alpha\leq\frac{5}{27}\approx 0.185$, and $k_{\alpha}\leq 4$ for $\alpha\leq\frac{57}{256}\approx 0.222$. In Figure 2, we plot the result we would get by computing $\delta(k)$ exactly, and compare with the estimate given in Theorem 1.
[…]
Figure 2: A blue point at $(k^{*},\alpha^{*})$ implies that every election has an $\alpha^{*}$-dominating set of at most $k^{*}$ candidates (i.e, $k_{\alpha}\leq k^{*}$ for $\alpha\leq\alpha^{*}$). These points are derived by evaluating $\delta(k)$ directly, and the red line compares to the estimate from Lemma 2 that $\delta(k)\approx\sqrt{\frac{\pi}{8k}}$.
Remark 2.
While Theorem 1 is stated as an existence result, it is not hard to make it algorithmic and find small approximately dominating sets in expected polynomial time. Firstly, since the maximal lottery is the equilibrium of a two-player $m\times m$ zero-sum game, it can be computed in polynomial time by solving a linear program. The proof samples $k$ times from the maximal lottery, and argues that the result is a $(\frac{1}{2}-\delta(k))$-dominating set with positive probability since the expectation of $Z:=\frac{1}{n}\sum_{v\in V}\delta_{v}$ is at most $\delta(k)$. We can instead use Markov’s inequality to argue that $Z$ is at most $(1+\eta)\delta(k)$ with probability at least $\frac{\eta}{1+\eta}$. Taking $\eta=\frac{1}{m}$ for example, by repeating the sampling process until it succeeds, we can find a $(\frac{1}{2}-(1+o(1))\delta(k))$-dominating set in expected polynomial time.
@@ Appendix A A note on lower bounds
We will argue that a random subset $B$ of $t]\setminus A$ of size $b$ satisfies this property with positive probability. For each $i$, the probability that $|A_{i}\cap B|>\frac{b}{2}$ is at most (very loosely)